OpenStudy (superhelp101):

Part 1 Create a difference of squares binomial or a perfect square trinomial that can be factored multiple times. Part 2 Provide the factors of this polynomial. Part 3 Explain, in complete sentences, the process you used to create the binomial/trinomial. @texaschic101

OpenStudy (superhelp101):

@Abhisar @iPwnBunnies @sweetburger

zepdrix (zepdrix):

Factored multiple times? How bout something like this?\[\Large\rm x^4-16\]=o

OpenStudy (superhelp101):

oh okay I didn't seem to understand "factored multiple times part"

zepdrix (zepdrix):

Or do they want difference of only squares, written as a binomial power?\[\Large\rm (x^2-1)^2\]

zepdrix (zepdrix):

written as a binomial to some power*

zepdrix (zepdrix):

That would be another option we could look at.

OpenStudy (superhelp101):

oh okay. is it possible to explain that also?

zepdrix (zepdrix):

So the way I'm starting this is by thinking of my difference of squares formula:\[\Large\rm a^2-b^2=(a-b)(a+b)\]So I know that if I can write both numbers as perfect squares (with subtraction between them) then they'll factor nicely using this identity.\[\Large\rm x^2-1=x^2-1^2\]Factoring to,\[\Large\rm x^2-1^2=(x-1)(x+1)\]Yes? Maybe if we stuck with the 16 it would be easier to work with.

OpenStudy (superhelp101):

Maybe I think the 16 will be easier

OpenStudy (superhelp101):

Don't you think?

zepdrix (zepdrix):

Fine fine fine :)\[\Large\rm x^2-16 \quad=\quad x^2-4^2\]Do you understand how we could use the formula from here?

OpenStudy (superhelp101):

yes so far. :)

zepdrix (zepdrix):

The reason I put the square on the outside of the brackets is because that's an easy way to get some factors!! c: Remember a square tells you to multiply the `thing` by itself.

zepdrix (zepdrix):

\[\Large\rm (x^2-16)^2=(x^2-16)(x^2-16)\]

OpenStudy (superhelp101):

oh okay would I factor that further?

OpenStudy (superhelp101):

since it's difference of two squares binomial

zepdrix (zepdrix):

Writing 16 as 4^2 shows us that it's a perfect square. \[\Large\rm (x^2-16)^2=(x^2-4^2)(x^2-4^2)\]

zepdrix (zepdrix):

Then we can use our difference of squares formula on each set.\[\Large\rm =(x-4)(x+4)(x-4)(x+4)\]

zepdrix (zepdrix):

I hope I didn't misread part 1.. The directions are a little strange :d

OpenStudy (superhelp101):

lol

OpenStudy (superhelp101):

thank you!

zepdrix (zepdrix):

Yah I hope that made some sense to you! :o